# Leveraging differentiable programming in the inverse problem of neutron stars

### v1

@@ II.5 Equation of state constraints
In this work, we consider constraints on the EOS from chiral effective field theory ($\chi$EFT), mass measurements of heavy pulsars, mass-radius constraints provided by the NICER experiment, and the tidal deformabilities measured in GW170817. Below, we provide more details on each of these constraints.
@@ II.5.1 Chiral effective field theory
At low energies, $\chi$EFT provides an expansion of the nuclear Hamiltonian in terms of the nucleon momenta over a breakdown scale Epelbaum et al. (2009); Machleidt and Entem (2011). After truncating the expansion at a certain order, the nuclear many-body problem involving the Hamiltonian can be solved numerically, from which the EOS follows. The truncation introduces theoretical uncertainty, which, however, can be systematically estimated Epelbaum et al. (2015); Drischler et al. (2020a, b); Armstrong et al. (2025). As a result, $\chi$EFT provides an uncertainty band for the EOS, bounded by two pressure curves $p_{-}(n)$ and $p_{+}(n)$. We constrain the EOS using this band, following an approach similar to Ref. Koehn et al. (2024a). In particular, the calculation done by Ref. Koehn et al. (2024a) yields a score function $f(p,n)$ that quantifies the consistency between a given EOS and the $\chi$EFT prediction. Given this score function, the likelihood is given by
$P(\bm{\theta}_{\rm{EOS}}|d_{\chi\rm{EFT}})\propto\exp\left(\int_{0.75n_{\rm{sat}}}^{n_{\rm{break}}}\frac{\log f(p(\bm{\theta}_{\rm{EOS}};n),n)}{n_{\rm{break}}-0.75n_{\rm{sat}}}\right)\,.$ (9)
We utilize the pressure band from Ref. Tews et al. (2018a), which extended calculations for pure neutron matter to $\beta$-equilibrium, making it applicable to NS EOS.
@@ II.5.2 Mass measurements of pulsars
Mass measurements of the heaviest known pulsars, obtained by evaluating relativistic effects in radio timing, provide a lower bound for the maximal mass of an NS, also referred to as the TOV mass $M_{\rm{TOV}}$. Here, we consider three massive pulsars: PSR J1614-2230 Demorest et al. (2010); Shamohammadi et al. (2023) with a mass of $1.937\pm 0.014\ M_{\odot}$, PSR J0348+0432 with a mass of $2.01\pm 0.04\ M_{\odot}$ Antoniadis et al. (2013) [^3] Note that recent work Saffer et al. (2024) argues the mass to be $\sim 1.8M_{\odot}$. However, in this work, we stick to the mass estimate determined previously, to allow for a direct comparison to Ref. Koehn et al. (2024a). and PSR J0740+6620 with a mass of $2.08\pm 0.07\ M_{\odot}$ (but see also Refs. Cromartie et al. (2019); Agazie et al. (2023)). All values are quoted at the $68\%$ credible level. Radio timing measurements $d_{\rm{radio}}$ give a posterior on the pulsar mass $P(M|d_{\rm{radio}})$, which is well approximated by a normal distribution with the mean and standard deviation values quoted above. For each pulsar, the likelihood on the EOS parameters is then
@@ II.5.3 NICER
Pulse profile modeling of NICER data has provided mass-radius posteriors for various pulsars Riley et al. (2019); Miller et al. (2019a); Riley et al. (2021); Miller et al. (2021); Choudhury et al. (2024); Salmi et al. (2024a). In this work, we limit ourselves to those of PSR J0030+0451 and PSR J0740+6620. These analyses have later been refined with more accumulated data and advanced models for the hotspots and instrument response Vinciguerra et al. (2024); Salmi et al. (2024b); Dittmann et al. (2024a). We use posterior samples on the masses and radii from the public data releases Riley et al. (2022); Miller et al. (2019b); Salmi et al. (2024c); Dittmann et al. (2024b). Their distributions $P(M,R|d_{\rm{NICER}})$ are approximated by a Gaussian kernel density estimate. In our inference, we sample the mass $M$ of each pulsar from a uniform prior distribution and compute the likelihood as
@@ II.5.4 GW170817
Finally, we consider constraints from tidal deformability measurements of GW170817.444The second observed BNS candidate so far, GW190425 Abbott et al. (2020), does not significantly constrain the EOS due to its lower signal-to-noise ratio and larger component masses, and therefore will not be considered in this work. For this, one first has to determine the posterior distributions on the source parameters $\bm{\theta}_{\rm{GW}}$ of GW170817 Veitch and Vecchio (2010); Veitch et al. (2015); Romero-Shaw et al. (2020). Here, we take the posterior samples produced by the analysis of Ref. Wouters et al. (2024), which obtained the posterior using Jim Edwards et al. (2023); Wong et al. (2023b); Wouters et al. (2024), a GW inference toolkit implemented in jax able to perform fast Bayesian inference on GW signals. The samples are publicly available at Ref. Wouters (2024).
From the posterior on the detector-frame chirp mass, the mass ratio, and the luminosity distance, we infer the source-frame component masses $M_{i}$ using the linear Hubble law, where $i=1,2$ labels the two components of the binary. For the inference on the EOS, we then consider the marginalized posterior distribution on $M_{i}$ and the tidal deformabilities $\Lambda_{i}$. Contrary to the case of the NICER mass-radius posteriors, a Gaussian kernel density estimate provides a poor approximation of this distribution. Therefore, we train [^5] Training this NF takes less than $\lesssim 15$ minutes on an NVIDIA A100 GPU using FlowJAX Ward (2024). an NF to obtain a tractable and more accurate approximation of the marginal posterior density, which we denote by $P_{\rm{NF}}(M_{i},\Lambda_{i}|d_{\rm{GW}})$. During inference, we sample the binary component masses $M_{i}$ and use the likelihood function
@@ II.6 Gradient-based inversion
In practice, we consider $N=500$ masses $M_{i}$, evenly spaced in the range $[1.0\ M_{\odot},M_{\rm{TOV}}]$, and the corresponding radii $\hat{R}_{i}$ and tidal deformabilities $\hat{\Lambda}_{i}$ of the “target” EOS, denoted by a hat. We have found that the results of the recovery do not significantly depend on the value of $N$ and the spacing of the masses. The lower bound on the mass is conservatively chosen below the lowest reliably observed NS mass to date Martinez et al. (2015)  [^6] Note that an NS candidate with a mass of $\sim 0.77M_{\odot}$ was proposed in Ref. Doroshenko et al. (2022), though.. We then measure the distance between a sample $\bm{\theta}_{\rm{EOS}}$, with corresponding values for the tidal deformabilities $\Lambda_{i}(\bm{\theta}_{\rm{EOS}})$ at $M_{i}$, and the target NS values with the following loss function:
@@
By starting from a randomized initial position in the EOS parameter space and iteratively performing gradient descent on this loss function, we aim to recover a set of EOS parameters that matches the given mass-tidal-deformability relation. In particular, we use the Adam optimizer, which stabilizes the descent with adaptive estimates of the first and second moments of the gradients Kingma and Ba (2014). Figure 1 illustrates our optimization algorithm. The NS curve of the EOS at the final iteration deviates from the target NS family with an error less than $100$ meters in radius and less than $10$ in tidal deformability for the mass range considered during optimization.
Figure 1: Illustration of the gradient-based inversion scheme recovering a given mass-radius relation of NSs. The color gradient shows the different iterations while performing gradient descent on the objective function of Eq. (13). The gradient is computed with respect to the EOS parameters throughout the TOV equations using automatic differentiation. The top panel illustrates the trajectory of the EOS parameters, showing only $L_{\rm{sym}}$ and $K_{\rm{sat}}$ for simplicity.
While preparing our manuscript, Ref. Lindblom and Zhou (2025) also introduced a recovery algorithm based on optimization. However, their method uses a spectral expansion Lindblom and Zhou (2024) and relatively low-dimensional EOS parameter spaces. Our scheme, instead, scales well to high-dimensional parameter spaces by employing gradient descent. Indeed, the illustration shown in Fig. 1 performed the inversion on the complete EOS parametrization described in Sec. II.1, using eight grid points for the CSE, and was able to recover the EOS. Since this parametrization involves $26$ parameters in total, this example demonstrates the applicability of our method in high-dimensional EOS spaces.
@@ III.1 Validation and computational scaling
In order to validate our method, we compare our results to those of Ref. Koehn et al. (2024a) which used an identical EOS parametrization. [^7] The quoted results for Ref. Koehn et al. (2024a) are obtained from https://multi-messenger.physik.uni-potsdam.de/eos_constraints/. Moreover, we use the same prior distribution, given in Table 1. We take eight grid points for the CSE parametrization.
@@
In Table 2, we compare the $95\%$ credible intervals of the inferred radius of a $1.4M_{\odot}$ NS, $R_{1.4}$. The first rows of the table correspond to inferences employing the constraints discussed in Sec. II.5 individually. The bottom row shows the result when considering all constraints jointly in a single Bayesian inference run. The table demonstrates that our results qualitatively agree with Ref. Koehn et al. (2024a). However, the present work directly samples the EOS parameters, whereas Ref. Koehn et al. (2024a) generated a fixed set of EOS candidates which were weighted using the constraints. This difference in sampling strategy can result in quantitative differences for the inferred $R_{1.4}$, as our approach allows for a higher effective sample size and more thorough exploration of the parameter space. Nevertheless, for all constraints considered here, we do observe quantitative agreement between both approaches for $M_{\rm{TOV}}$, the pressure of the EOS at $3n_{\rm{sat}}$, and the maximal central density inside a NS, $n_{\rm{TOV}}$, as shown by Table 3 in Appendix A.
Table 2: Comparison of inferred $R_{1.4}$ ($95\%$ credible interval) in kilometers between Ref. Koehn et al. (2024a) and this work for different EOS constraints. The bottom row refers to the posterior using all listed constraints jointly.
@@
Constraint Ref. Koehn et al. (2024a) This work
@@
After validating that our framework agrees with previously published results, we now assess the runtime and its scaling as a function of the dimensionality of the EOS parameter space. Using a single NVIDIA H100 GPU, the inferences listed in Table 2 were performed under $1$ hour of wall time. Solving the TOV equations to obtain the macroscopic properties of NSs ($100$ masses and their corresponding radii, and tidal deformabilities) for a single proposal takes around $\sim 0.3$ milliseconds, which is on par with previous works relying on machine learning surrogates (see, e.g., Refs Reed et al. (2024); Magnall et al. (2024); Tiwari and Pai (2024) for recent examples using neural networks), however, without the need of pretraining a machine learning model in our case.
Future experiments are expected to provide more precise observations of NS, which could require a more flexible EOS parameterization to avoid biases. This motivates us to extend the parameter space by adding more CSE grid points and assess the impact on the scaling of our sampling efficiency. Specifically, we perform Bayesian inference jointly on all EOS constraints considered in this work by testing CSE grids with $10$, $20$, and $30$ points, respectively. Figure 2 shows the runtime per effective sample size (ESS), estimated with Geyer’s initial monotone sequence criterion Geyer (2011, 1992). The axis on the right converts this number to an estimate of the total wall time required in case one requires $5000$ effective samples. We perform the scaling runs on two different GPU architectures, namely an NVIDIA A100 and NVIDIA H100, and show the average and standard deviation over all the EOS parameters. The parametrization with the highest number of grid points increases the runtime roughly by a factor $4$ to produce a similar number of effective samples. Therefore, our inferences still only require a couple of hours to provide accurate posteriors on the EOS, even when using a more flexible parametrization.
@@ III.2 Constraining $n_{\rm{break}}$ with neutron star data
The EOS parametrization employed in this work takes the breakdown density $n_{\rm{break}}$ of the metamodel as a parameter that can be varied freely. This density is an indicator for the densities for which a simple nucleonic description of the EOS is not sufficient anymore. We can use our framework to constrain $n_{\rm{break}}$ using NS data and directly probe this breakdown with astrophysical information. To this end, we perform Bayesian inference as explained above with two key modifications. First, we use a wider, agnostic prior on $n_{\rm{break}}$, namely a uniform distribution between $[1n_{\rm{sat}},4n_{\rm{sat}}]$. Second, we only consider EOS constraints involving NS data, i.e., we do not use the constraint from $\chi_{\rm{EFT}}$. Figure 3 shows the marginal posterior distribution of $n_{\rm{break}}$. We infer a breakdown density of $n_{\rm{break}}=2.47_{-1.02}^{+1.53}\ n_{\rm{sat}}$ at the $90\%$ credible level. Hence, NS data prefers a smooth nucleonic EOS up to 2-3 times the nuclear saturation density, a range within the reach of microscopic nuclear theory Essick et al. (2020a). While $n_{\rm{break}}$ is not tightly constrained by the data considered in this work, future experiments providing more NS data can potentially improve its precision. Nevertheless, our result serves as proof of concept that the metamodel breakdown density can be determined on-the-fly with Bayesian inference.
We point out that Ref. Biswas et al. (2021) similarly inferred this transition density. However, comparing the two results is complicated, as they use different parameterizations and prior choices. In particular, we sample the $n_{\rm{break}}$ parameter on-the-fly during inference, whereas Ref. Biswas et al. (2021) infers this parameter by considering Bayes factors of different runs with fixed $n_{\rm{break}}$. Moreover, we extend the MM parametrization to higher orders and consider a speed-of-sound parametrization at high densities, rather than piecewise polytropes.
Figure 3: Posterior on $n_{\rm{break}}$ inferred from a uniform prior between $1n_{\rm{sat}}$ to $4n_{\rm{sat}}$ (other priors are given in Table 1) and given the NS observations from Table 2.

### v2

@@ II.5 Equation of state constraints
[…]
@@ II.5.1 Chiral effective field theory
[…]
$P(\bm{\theta}_{\rm{EOS}}|d_{\chi\rm{EFT}})\propto\exp\left(\int_{0.75n_{\rm{sat}}}^{n_{\rm{break}}}\frac{\log f(p(\bm{\theta}_{\rm{EOS}};n),n)}{n_{\rm{break}}-0.75n_{\rm{sat}}}\mathop{}\!\mathrm{d}n\right)\,.$ (9)
We utilize the pressure band from Ref. Tews:2018kmu, which extended calculations for pure neutron matter to $\beta$-equilibrium, making it applicable to NS EOS.
The approach used in this work neglects the correlations between pressures. However, we follow the same approach as Ref. Koehn:2024set in order to fully match their setup for our comparison in Sec. III.1.
@@ II.5.2 Mass measurements of pulsars
Mass measurements of the heaviest known pulsars, obtained by evaluating relativistic effects in radio timing, provide a lower bound for the maximal mass of an NS, also referred to as the TOV mass $M_{\rm{TOV}}$. Here, we consider three massive pulsars: PSR J1614-2230 Demorest:2010bx; Shamohammadi:2022ttx with a mass of $1.937\pm 0.014\ M_{\odot}$, PSR J0348+0432 with a mass of $2.01\pm 0.04\ M_{\odot}$ Antoniadis:2013pzd [^3] Note that recent work Saffer:2024tlb argues the mass to be $\sim 1.8M_{\odot}$. However, in this work, we stick to the mass estimate determined previously, to allow for a direct comparison to Ref. Koehn:2024set. and PSR J0740+6620 with a mass of $2.08\pm 0.07\ M_{\odot}$ (but see also Refs. NANOGrav:2019jur; NANOGrav:2023hde). All values are quoted at the $68\%$ credible level. Radio timing measurements $d_{\rm{radio}}$ give a posterior on the pulsar mass $P(M|d_{\rm{radio}})$, which is well approximated by a normal distribution with the mean and standard deviation values quoted above. For each pulsar, the likelihood on the EOS parameters is then
@@ II.5.3 NICER
Pulse profile modeling of NICER data has provided mass-radius posteriors for various pulsars Riley:2019yda; Miller:2019cac; Riley:2021pdl; Miller:2021qha; Choudhury:2024xbk; Salmi:2024bss. In this work, we limit ourselves to those of PSR J0030+0451 and PSR J0740+6620. These analyses have later been refined with more accumulated data and advanced models for the hotspots and instrument response Vinciguerra:2023qxq; Salmi:2024aum; Dittmann:2024mbo. We use posterior samples on the masses and radii from the public data releases NICER_posterior_samples_Zenodo_J0030_Riley; NICER_posterior_samples_Zenodo_J0030_Miller; NICER_posterior_samples_Zenodo_J0740_Salmi; NICER_posterior_samples_Zenodo_J0740_Dittmann. Their distributions $P(M,R|d_{\rm{NICER}})$ are approximated by a Gaussian kernel density estimate. In our inference, we sample the mass $M$ of each pulsar from a uniform prior distribution and compute the likelihood as
@@ II.5.4 GW170817
Finally, we consider constraints from tidal deformability measurements of GW170817.444The second observed BNS candidate so far, GW190425 LIGOScientific:2020aai, does not significantly constrain the EOS due to its lower signal-to-noise ratio and larger component masses, and therefore will not be considered in this work. For this, one first has to determine the posterior distributions on the source parameters $\bm{\theta}_{\rm{GW}}$ of GW170817 Veitch:2009hd; Veitch:2014wba; Romero-Shaw:2020owr. Here, we take the posterior samples produced by the analysis of Ref. Wouters:2024oxj, which obtained the posterior using Jim Edwards:2023sak; Wong:2023lgb; Wouters:2024oxj, a GW inference toolkit implemented in jax able to perform fast Bayesian inference on GW signals. The samples are publicly available at Ref. Jim_GW170817_posterior_samples.
From the posterior on the detector-frame chirp mass, the mass ratio, and the luminosity distance, we infer the source-frame component masses $M_{i}$ using the linear Hubble law, where $i=1,2$ labels the two components of the binary. For the inference on the EOS, we then consider the marginalized posterior distribution on $M_{i}$ and the tidal deformabilities $\Lambda_{i}$. Contrary to the case of the NICER mass-radius posteriors, a Gaussian kernel density estimate provides a poor approximation of this distribution. Therefore, we train [^5] Training this NF takes less than $\lesssim 15$ minutes on an NVIDIA A100 GPU using FlowJAX ward2023flowjax. an NF to obtain a tractable and more accurate approximation of the marginal posterior density, which we denote by $P_{\rm{NF}}(M_{i},\Lambda_{i}|d_{\rm{GW}})$. During inference, we sample the binary component masses $M_{i}$ and use the likelihood function
@@ II.6 Gradient-based inversion
[…]
@@
We note that the loss function is based on mock data coming from the “true” EOS rather than existing NS observations, and therefore is introduced in this work as a proof of concept rather than a full replacement of Bayesian inference.
By starting from a randomized initial position in the EOS parameter space and iteratively performing gradient descent on this loss function, as shown in Eq. (7), we aim to recover a set of EOS parameters that matches the given mass-tidal-deformability relation. The derivative of the loss function is computed with automatic differentiation, as explained in Sec. II.3. In particular, we use the Adam optimizer, which stabilizes the descent with adaptive estimates of the first and second moments of the gradients Kingma:2014vow. Figure 1 illustrates our optimization algorithm. The NS curve of the EOS at the final iteration deviates from the target NS family with an error less than $100$ meters in radius and less than $10$ in tidal deformability for the mass range considered during optimization. We note that we show the optimization run after truncating around $600$ iteration steps for visualization purposes. We expect further improvement in the convergence, especially at lower masses, if the optimizer runs for more iterations. Since the “true” EOS considered in this example comes from the same EOS parametrization as used in the recovery, a perfect recovery of the EOS is theoretically possible. However, in practice, the optimization can get stuck in local minima. This might be more problematic in case the EOS does not follow the same parametric form as used in the recovery. Moreover, incorporating real data with uncertainties will change the loss function of Eq. (13) and the robustness of the recovery procedure. These additional complications will be addressed in future work.
Figure 1: Illustration of the gradient-based inversion scheme recovering a given mass-radius relation of NSs. The color gradient shows the different iterations while performing gradient descent on the objective function of Eq. (13). The gradient is computed with respect to the EOS parameters throughout the TOV equations using automatic differentiation. Top panel: Trajectory of the EOS parameters, showing only $L_{\rm{sym}}$ and $K_{\rm{sat}}$ for simplicity. Middle panel: Mass-radius and mass-tidal deformability curves during the gradient descent, with the target shown in black. Lower panel: Pressure as a function of density curves during optimization.
While preparing our manuscript, Refs. Lindblom:2025kjz; Li:2025obt also introduced a recovery algorithm based on optimization. However, their methods differ from ours in several ways. Ref. Lindblom:2025kjz uses a spectral expansion Lindblom:2024yio and relatively low-dimensional EOS parameter spaces. Our scheme, instead, scales well to high-dimensional parameter spaces by employing gradient descent. Indeed, the illustration shown in Fig. 1 performed the inversion on the complete EOS parametrization described in Sec. II.1, using eight grid points for the CSE, and was able to recover the EOS. Since this parametrization involves $26$ parameters in total, this example demonstrates the applicability of our method in high-dimensional EOS spaces. Ref. Li:2025obt uses a gradient-based inversion scheme, but uses a neural network as EOS parametrization and relies on linear response analysis to compute the gradients. In contrast, we use automatic differentiation, which enables us to compute these gradients for any EOS parametrization.
@@ III.1 Validation and computational scaling
In order to validate our method, we compare our results to those of Ref. Koehn:2024set which used an identical EOS parametrization. [^7] The quoted results for Ref. Koehn:2024set are obtained from https://multi-messenger.physik.uni-potsdam.de/eos_constraints/. Moreover, we use the same prior distribution, given in Table 1. We take eight grid points for the CSE parametrization.
@@
In Table 2, we compare the $95\%$ credible intervals of the inferred radius of a $1.4M_{\odot}$ NS, $R_{1.4}$. The first rows of the table correspond to inferences employing the constraints discussed in Sec. II.5 individually. The bottom row shows the result when considering all constraints jointly in a single Bayesian inference run. The table demonstrates that our results qualitatively agree with Ref. Koehn:2024set, but show some quantitative differences in the inferred radii. Nevertheless, for all constraints considered here, we do observe quantitative agreement between both approaches for $M_{\rm{TOV}}$, the pressure of the EOS at $3n_{\rm{sat}}$, and the maximal central density inside a NS, $n_{\rm{TOV}}$, as shown by Table 3 in Appendix A.
[…]
Table 2: Comparison of inferred $R_{1.4}$ ($95\%$ credible interval) in kilometers between Ref. Koehn:2024set and this work for different EOS constraints. The bottom row refers to the posterior using all listed constraints jointly.
@@
Constraint Ref. Koehn:2024set This work
@@
After validating that our framework agrees with previously published results, we now assess the runtime and its scaling as a function of the dimensionality of the EOS parameter space. Using a single NVIDIA H100 GPU, the inferences listed in Table 2 were performed under $1$ hour of wall time. Solving the TOV equations to obtain the macroscopic properties of NSs ($100$ masses and their corresponding radii, and tidal deformabilities) for a single proposal takes around $\sim 0.3$ milliseconds, which is on par with previous works relying on machine learning surrogates (see, e.g., Refs Reed:2024urq; Magnall:2024ffd; Tiwari:2024jui for recent examples using neural networks), however, without the need of pretraining a machine learning model in our case.
In the future, more precise observations of NSs will become available, which could require a more flexible EOS parameterization to avoid biases. This motivates us to extend the parameter space by adding more CSE grid points and assess the impact on the scaling of our sampling efficiency. Specifically, we perform Bayesian inference jointly on all EOS constraints considered in this work by testing CSE grids with $10$, 2020, and $30$ points, respectively. Figure 2 shows the runtime per effective sample size (ESS), estimated with Geyer’s initial monotone sequence criterion geyer_SSE_ref1; geyer_SSE_ref2. The axis on the right converts this number to an estimate of the total wall time required in case one requires $5000$ effective samples. We perform the scaling runs on two different GPU architectures, namely an NVIDIA A100 and NVIDIA H100, and show the average and standard deviation over all the EOS parameters. The parametrization with the highest number of grid points increases the runtime roughly by a factor $4$ to produce a similar number of effective samples. Therefore, our inferences still only require a couple of hours to provide accurate posteriors on the EOS, even when using a more flexible parametrization.
@@ III.2 Constraining $n_{\rm{break}}$ with neutron star data
The EOS parametrization employed in this work takes the breakdown density $n_{\rm{break}}$ of the metamodel as a parameter that can be varied freely, after which the CSE part of the parametrization is used for the high-density EOS. This density is an indicator of the densities for which a simple nucleonic description of the EOS is not sufficient anymore. We can use our framework to constrain $n_{\rm{break}}$ using NS data and directly probe this breakdown with astrophysical information. To this end, we perform Bayesian inference as explained above with two key modifications. First, we use a wider, agnostic prior on $n_{\rm{break}}$, namely a uniform distribution between $1n_{\rm{sat}},4n_{\rm{sat}}]$. The priors on the other MM and CSE parameters are identical to those shown in Tab. 1. Second, we only consider EOS constraints involving NS data, i.e., we do not use the constraint from $\chi_{\rm{EFT}}$. Figure 3 shows the marginal posterior distribution of $n_{\rm{break}}$. We infer a breakdown density of $n_{\rm{break}}=2.47_{-1.02}^{+1.53}\ n_{\rm{sat}}$ at the $90\%$ credible level. Hence, NS data prefers a smooth nucleonic EOS up to 2-3 times the nuclear saturation density, a range within the reach of microscopic nuclear theory Essick:2020flb. While $n_{\rm{break}}$ is not tightly constrained by the data considered in this work, future detectors providing more NS data can potentially improve its precision. Nevertheless, our result serves as proof of concept that the metamodel breakdown density can be determined on-the-fly with Bayesian inference.
We point out that Ref. Biswas:2020puz similarly inferred this transition density. However, comparing the two results is complicated, as they use different parameterizations and prior choices. In particular, we sample the $n_{\rm{break}}$ parameter on the fly during inference, whereas Ref. Biswas:2020puz infers this parameter by considering Bayes factors of different runs with fixed $n_{\rm{break}}$. Moreover, we extend the MM parametrization to higher orders and consider a speed-of-sound parametrization at high densities, rather than piecewise polytropes.
Figure 3: Posterior on $n_{\rm{break}}$ of the metamodel parametrization inferred from a uniform prior between $1n_{\rm{sat}}$ to $4n_{\rm{sat}}$ (other priors are given in Tab. 1) and given the NS observations from Tab. 2.